Use the intercepts to graph the equation.
x+3y=6
(graph is 10 to -10 by 2)

Answer by Fombitz(32388)

When x=0,
x+3y=6
0+3y=6
y=2
The y intercept is (0,2).
.
.
.
When y=0,
x+3y=6
x+0=6
x=6
The x=intercept is (6,0).
.
.
.
Graph those two points.

.
.
.
Now draw the line connecting those two points and extend the line to the edges of your graph.
.
.
.